OpenStudy (anonymous):

if f(x)=summation of x^4n/ (4n)! prove that the fourth derivative of f(x)= f(x)

OpenStudy (anonymous):

summation also known as sigma

ganeshie8 (ganeshie8):

*

OpenStudy (ybarrap):

$$ \Large{ f(x)=\Sigma {x^{4n}\over{(4n)!}}\\ =\cfrac{1}{1}+\cfrac{x^8}{8!}+\cfrac{x^{12}}{12!}+\cfrac{x^{16}{}}{16!}+\cdots } $$ Is this the problem? Now find \(\large{\cfrac{\partial^4 f(x)}{\partial x^4}}\) and compare to f(x) Note that $$ \large{ f(x)=\Sigma {x^{4n}\over{(4n)!}}\\ = \cfrac{1}{2} (\cos(x)+\cosh(x)-2) } $$ which might be easier to differential 4 times.

OpenStudy (ybarrap):

Correction on expansion: $$ \Large{ f(x)=\Sigma {x^{4n}\over{(4n)!}}\\ =\cfrac{1}{1}+\cfrac{x^4}{4!}+\cfrac{x^{8}}{8!}+\cfrac{x^{12}{}}{12!}+\cdots } $$

OpenStudy (anonymous):

yes that is the correct problem. so i do the expansion?

OpenStudy (anonymous):

and then how do you get the 1/2 (cosx +coshx -2?

OpenStudy (ybarrap):

I used a table of series expansions, which is usually a legitimate method if you stumble upon it.

OpenStudy (anonymous):

oh so from there you find that its 1/2(cosx+coshx -2) ? what does the h stand for though?

OpenStudy (anonymous):

Another way, done by differentiating the power series: \[\large\begin{align*}f(x)&=\sum_{n=0}^\infty\frac{x^{4n}}{(4n)!}\\\\ f'(x)&=\sum_{n=0}^\infty \frac{4nx^{4n-1}}{(4n)!}\\ &=\sum_{n=1}^\infty \frac{x^{4n-1}}{(4n-1)!}\\\\ f''(x)&=\sum_{n=1}^\infty \frac{(4n-1)x^{4n-2}}{(4n-1)!}\\ &=\sum_{n=1}^\infty \frac{x^{4n-2}}{(4n-2)!}\\\\ f'''(x)&=\sum_{n=1}^\infty\frac{(4n-2)x^{4n-3}}{(4n-2)!}\\ &=\sum_{n=1}^\infty\frac{x^{4n-3}}{(4n-3)!}\\\\ f^{(4)}(x)&=\sum_{n=1}^\infty\frac{(4n-3)x^{4n-4}}{(4n-3)!}\\ &=\sum_{n=1}^\infty\frac{x^{4n-4}}{(4n-4)!} \end{align*}\] Then check each sum term by term.

OpenStudy (anonymous):

Wolfram insists that the starting index is \(n=0\) though...

OpenStudy (ybarrap):

cosh(x) is cosine hyperbolic its derivative is sinh(x) and derivative of sinh(x) is cosh(x)

OpenStudy (anonymous):

@ssoleman, \[\cosh x=\frac{e^x+e^{-x}}{2}~~\text{and}~~\sinh x=\frac{e^x-e^{-x}}{2}\] are the definitions for the hyperbolic trig functions.

OpenStudy (anonymous):

thank you sithsandgiggles.. ill look into that way too but so the derivative of cosh(x) is not -sinh(x)?

OpenStudy (ybarrap):

no, d/dx cosh(x) = sinh(x)

OpenStudy (anonymous):

\[\frac{d}{dx}\cosh x=\frac{d}{dx}\left[\frac{e^x+e^{-x}}{2}\right]=\frac{e^{x}-e^{-x}}{2}=\sinh x\]

OpenStudy (anonymous):

okay thank you.. and now that i have found all the derivative until the fourth, where do I go from there with this ?

OpenStudy (ybarrap):

now compare to f(x)

OpenStudy (anonymous):

so does it make sense that the fourth d of f(x) = 1/2(cos(x)+cosh(x)

OpenStudy (anonymous):

which means that they equal eachother!